OpenStudy (anonymous):

3√18 + 2√72 simplify

OpenStudy (anonymous):

\[21\sqrt{2}\]

OpenStudy (anonymous):

can you do the work plz?

OpenStudy (anonymous):

\[3\sqrt{18}=9\sqrt{2} 2\sqrt{72}=12\sqrt{2} 9\sqrt{2}+12\sqrt{2}=21\sqrt{2}\]

OpenStudy (anonymous):

sorry it got messed up im new to this site

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

it is 3root18 = 9root2 2root72 = 12root2 12root2 + 9root2= 21root2